OpenStudy (anonymous):

If x and y are integers, and 3x + 2y = 13, which of the following could be the value of y ? Select ALL values that apply. B. 1 C. 2 D. 3

OpenStudy (studygurl14):

is there any info about the value of x?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

i think its C

OpenStudy (anonymous):

2x2 is 4. 13-3 is 9 therefore x would be 3

OpenStudy (anonymous):

13-4*

OpenStudy (anonymous):

if you solve for the other possible values x would not end up being an integer

OpenStudy (studygurl14):

3x + 2y = 13 3x + 2(1) = 13 3x + 2 = 13 3x = 11 x = 11/3 Not an integer 3x + 2y = 13 3x + 2(2) = 13 3x + 4 = 13 3x = 9 x = 3 An integer 3x + 2y = 13 3x + 2(3) = 13 3x + 6 = 13 3x = 7 x = 7/3 Not an integer It is C.